### v1

# From #Dr00gtiktok to #harmreduction: Exploring Substance Use Hashtags on TikTok

Abstract
The rise of TikTok as a primary source of information for youth, combined with its unique short-form video format, creates urgent questions about how substance use content manifests and spreads on the platform. This paper provides the first in-depth exploration of substance use-related content on TikTok, covering all major substance categories as classified by the Drug Enforcement Agency. Through social network analysis and qualitative coding, we examined more than 2,333 hashtags across 39,509 videos, identified 16 distinct hashtag communities and analyzed their interconnections and thematic content. Our analysis revealed a highly interconnected small-world network where recovery-focused hashtags like “#addiction”, “#recovery”, and “#sober” serve as central bridges between communities. Through manual coding of 351 representative videos, we found that Recovery Advocacy content (33.9%) and Satirical content (28.2%) dominate, while direct substance depiction appears in only 26% of videos, with active use shown in just 6.5% of them. This suggests TikTok functions primarily as a recovery support platform rather than a space promoting substance use. We found strong alignment between hashtag communities and video content, indicating organic community formation rather than attempts to evade content moderation. Our findings inform how platforms can balance content moderation with preserving valuable recovery support communities, while also providing insights for the design of social media-based recovery interventions.
Introduction
Social media platforms have become the primary lens through which young people view and understand the world, with TikTok emerging as a dominant force among youth - reaching 63% of teens, with 17% reporting near-constant use (Vogels, Gelles-Watnick, and Anderson 2023). More concerning is that TikTok has become a primary news source for this demographic, often replacing Google for information-seeking (Walker and Matsa 2023).
Figure 1: Examples of TikTok videos that can be found in the substance use communities based on hashtag use
This shift in information consumption patterns takes on particular urgency when considering content related to illicit substances. While overall drug use among 8th, 10th, and 12th graders showed promising declines during COVID-19 social distancing (National Institute on Drug Abuse 2022), this drop in usage may have masked how drug-related discussions were shifting online. Social media platforms can rapidly disseminate trends and behaviors through psychological processes like social comparison and fear of missing out (FOMO) (Montag, Yang, and Elhai 2021), with adolescents often imitating influencers and peers they follow. Although most social media platforms use content moderation, their rapid and viral nature can make it challenging to effectively regulate exposure to harmful content (Boyd 2014; Lookingbill and Le 2024), potentially amplifying dangerous trends. Therefore, understanding such emerging trends on social media platforms such as TikTok, Instagram, and Reddit (Rutherford et al. 2022a, 2023; Tan and Weinreich 2021; Bouzoubaa, Young, and Rezapour 2024) becomes more critical. For instance, the “xandemic” between 2016-2018, characterized by high school students documenting the recreational use of Xanax to the point of blackout, exemplifies how social media trends can blur the lines between casual and harmful behaviors and even normalize dangerous substance use behaviors (Fuller et al. 2024). Similarly, platform-specific hashtag “#pingtok” - a portmanteau of the slang term for an MDMA pill, pinger, and TikTok - which has accumulated over 26,000 posts at the time of writing, demonstrates how substance-related content can become embedded in platform culture, with its own specific terminology and community. These trends are especially concerning given robust evidence that early exposure to substance use significantly increases the risk of developing substance use disorders (SUD) in adulthood (McCabe et al. 2022; Volkow and Wargo 2022; Odgers et al. 2008).
Previous research on substance-related content on TikTok has typically focused on either specific substances such as cannabis or nicotine (Rutherford et al. 2023, 2022a; Morales, Fahrion, and Watkins 2022; Tan and Weinreich 2021) or isolated trending phenomenon like “#pingtok” (Whelan, Noller, and Ward 2024). While some researchers have examined hashtag networks and content moderation strategies around other sensitive topics like eating disorders (Bickham et al. 2024) and how hashtags create communities around shared experiences (Eriksson Krutrök 2021), there remains limited understanding of how hashtags shape substance-related discourse on TikTok. Moreover, no studies have comprehensively examined how substance-related content manifests across TikTok’s ecosystem of communities, particularly through its short-form video format and hashtag-driven discovery system (Abidin 2020). To address this gap, we draw from the Drug Enforcement Agency’s comprehensive drug classification system and examine content from every major substance category using a mixed-methods investigation of TikTok content. Our research questions are:
  • RQ1: What is the network structure of substance-related hashtags on TikTok, and what distinct communities emerge within this network?
  • RQ2: What are the topics and key characteristics of the videos associated with the hashtags in these communities?
  • RQ3: What are the effects of substance visibility on topic distribution and engagement metrics?
For this purpose, we developed a list of 585 seed hashtags related to substance use and collected more than a thousand videos for each of them, resulting in a dataset of 504,294 hashtags and 272,910 videos  [^1] Datasets, including seed hashtags, and code are available at https://github.com/social-nlp-lab/explore˙tiktok˙su. Videos are not directly provided but can be accessed via URLs.. After filtering relevant hashtags, we identified 2,333 exemplary hashtags covering the broad categories of substance use, and 351 videos for qualitative coding. Using a combination of network analysis of hashtag co-occurrence and in-depth content analysis, our analysis showed a highly interconnected small-world network where recovery-focused hashtags like “#addiction”, “#recovery”, and “#sober” serve as central bridges between communities. Within these communities, we found that Recovery Advocacy content (33.9%) and Satirical content (28.2%) dominate, while direct substance depiction appeared in only 26% of videos, with active use shown in just 6.5%. This suggests TikTok functions primarily as a recovery support platform rather than a space promoting substance use.
Through this work, we contribute to social computing and substance use literature by providing the first comprehensive mapping of substance-related content on TikTok. Our findings can inform platform policies around content moderation strategies that preserve supportive recovery communities while better identifying potentially harmful content, guide the development of social media-based recovery support interventions, and help design youth substance use prevention efforts that account for how substance-related content actually manifests on TikTok.
Related Works
As digital platforms become ubiquitous, understanding their impact, especially on young people’s health behaviors regarding substance use, is more crucial than ever. On the one hand, social media emerged as a medium for facilitating sensitive discussions and disclosure of lived experiences, including mental health and eating disorders (Rutherford et al. 2022a; Bickham et al. 2024; Bouzoubaa et al. 2024). On the other hand, several studies have examined the relationship between online peer groups, social media homophily, and substance use, finding positive correlations between substance use and social media engagement. For example, Willoughby et al. (2024) found that young adults reported an increased intention to use cannabis after exposure to pro-cannabis messages online. Similarly, Sun et al. (2023) showed that exposure to e-cigarette advertisements on social media platforms is associated with adolescents’ e-cigarette usage in later years. Miller et al. (2021) examined how online and offline peer interactions influence substance use behaviors and found that online belonging and social media homophily predicted regular stimulant and opioid use among substance users. Hanson et al. (2013) have studied online discussions related to stimulant drug use on Twitter and found a high correlation between the popularity of such topics and final exam periods. Geusens et al. (2023) found that higher levels of alcohol and marijuana consumption predicted later social media posts about those substances, but posting did not predict later substance use. Sun et al. (2023) found a positive correlation between early exposure to e-cigarette advertisement and subsequent lifetime usage.
Hashtags, a medium for interactions on platforms like Facebook or Twitter, have been studied comprehensively in the past (Mousavi and Ouyang 2021; Ince, Rojas, and Davis 2017; Small 2011; Rezapour et al. 2017). People can contribute to the narrative by hash-tagging their own personal experiences and thoughts (Yang 2016). At their core, hashtags facilitate content discovery and community creation by allowing users to contribute to the discussion through keyword tagging (Small 2011). Previous studies have examined substance use discourse across various social media platforms, contributing to our understanding of how substance use behaviors manifest online. For instance, on Instagram, investigating “#studydrugs” demonstrated a predominantly positive portrayal of stimulant use for academic performance enhancement (Petersen et al. 2021).
On TikTok, content creators deliberately employ hashtags like “#fyp” (For You Page) to increase visibility and maximize audience reach within TikTok’s recommendation system (Abidin 2020). The platform’s ranking algorithms connect users based on hashtag engagement patterns, creating what Eriksson Krutrök (2021) calls “algorithmic closeness” - subcommunities united through shared content interests. This algorithmic dynamic shapes how substance-related discourse spreads and evolves on TikTok. For example, research examining TikTok hashtags like “#puffbar” has revealed concerning patterns of youth-oriented substance marketing and normalization (Tan and Weinreich 2021), while analysis of “#stonertok” and “#420vibes” showed the majority of cannabis-related videos positively portray substances (Rutherford et al. 2022b).
Our study advances existing research by employing a network-based and qualitative approach to explore the interactions and communities associated with substance-related hashtags on TikTok. While previous research has thoroughly documented individual communities, the dynamics of how these groups interact across the platform’s broader ecosystem remain understudied. Our work can help social media platforms like TikTok refine their policies to better distinguish between harmful content requiring removal and harm reduction information that should be preserved, contributing to more effective content moderation.
Methods
Data
Seed Hashtag Selection: To develop an initial set of seed hashtags related to substance use for the data collection, we first identified hashtags previously used in studies of substance-related content on social media platforms (Petersen et al. 2021; Russell et al. 2022; Rutherford et al. 2023, 2022b; Singh and Wu 2021) as well as searching for hashtags on the platform associated with substance names (both formal and slang) listed on the Drug Enforcement Agency’s Fact Sheet (United States Drug Enforcement Administration 2025). Using this set, we conducted iterative searches on TikTok to identify additional relevant hashtags by examining co-occurring hashtags within the content. This process continued until we reached saturation, resulting in a final set of 585 substance-use-related hashtags. [^2] The list will be shared upon the paper’s acceptance.
Data Collection: Using the Unofficial TikTok API [^3] https://github.com/davidteather/TikTok-Api, we collected data for around 1,000 algorithmically served videos for each of the 585 seed hashtags. This resulted in a dataset of 275,291 videos and a total of 504,294 unique hashtags.
Figure 2: Data selection and pre-processing pipeline to create the filtered set.
Data Preprocessing and Network Refinement: To enhance the feasibility of our analysis and reduce noise from potentially irrelevant content, we filtered the data to include only videos that featured at least two of the seed hashtags. This filtering step resulted in a refined dataset of 39,602 videos and 99,727 unique hashtags (Figure 2). To further clean the data, we constructed a co-occurrence network where nodes represent individual hashtags and edges represent their co-occurrence in videos. More specifically, edge weights were defined by the frequency with which two hashtags appeared together in videos. We then employed a two-stage filtering process to focus on the most relevant and interconnected hashtags. First, we removed hashtags with low connectivity by filtering out nodes in the bottom 25th percentile of degree centrality. This step eliminated peripheral hashtags with minimal connections to the broader substance use discourse, removing those that were incidentally rather than substantively related to substance use discussions. Next, we applied k-core decomposition ($k=10$) to identify the densely connected core of the network (Seidman and Foster 1978). K-core decomposition recursively removes nodes with degrees less than $k$ until all remaining nodes have at least $k$ connections, ensuring a cohesive network structure. This threshold was chosen based on exploratory analysis, which revealed that $k=10$ produced the most structurally cohesive subset of hashtags while retaining a sufficient number of nodes and edges to analyze the discourse. The resulting refined network of hashtags contained 2,333 nodes and 46,990 edges, representing the most interconnected hashtags in substance use-related content.
Communities Definition Example Hashtags
Emotions and Feelings Words and phrases related to emotional states and feelings. #love, #gratitude
Health Conditions Words and phrases related to health issues directly related to substance use, addiction-related health problems, or chronic conditions that may lead to substance use. #addiction, #chronicpain
Alcohol Words related to alcoholic beverages and spirits. This is just about the substances or drinks and not the side effects and consumption methods. #alcohol, #vodka
Cannabis Substances related to marijuana, including recreational and medicinal use. #cannabis, #stonersoftiktok
Cognitive Enhancement Substances about nootropics, smart drugs, and methods to improve cognitive function. #modafinil, #smartdrugs
Commonly-Misused Substances Substances frequently used and misused, such as licit and illicit substances. #heroin, #xans
Consumption Method Words and phrases that show specific ways in which substances (including alcohol, illicit or licit drugs, tobacco, and cannabis) are ingested, administered, or used. #smoking, #injection
Awareness and Advocacy Words related to information and strategies to prevent and raise awareness of substance abuse or reduce associated harm, discussions about drug laws, policies, and related social issues, and content about addiction recovery, sobriety, and support systems. #harmreduction, #recoverylife
Other Substances Mentions of less common substances, oftentimes legal, over-the-counter medications, herbal remedies, or supplements. #vitamins, #magnesium
Platform Tags and features specific to social media engagement, visibility, and trending tactics. #fyp, #viral
Substance Effects Words and phrases that describe the physical or mental effects of substance or alcohol use, both desired effects and side effects. #stoned, #drunk
Tobacco Nicotine Words related to tobacco products, cigarettes, vaping, and nicotine use. #vaping, #ecigs
Humor Words, phrases, or hashtags related to jokes, memes, or any content meant to be funny but specific to substance use, addiction, or recovery. #soberhumor, #drughumor
Location Words related to geographical locations, including cities, states, countries, or continents. #kensingtonphilly, #boston
Occupation Words related to occupations or professions. #nurselife, #medstudent
Identity and Community Hashtags related to any social identity, demographic group, or community affiliation, such as race, ethnicity, gender identity, sexual orientation, disability status, socioeconomic background, immigration status, religion, age group, or membership in specific subcultures or communities. #lgbtqia, #transgender
Misc Any tag that does not fit into the above categories. #cookies, #foundersday
Table 1: Categories for semantically classifying hashtags from TikTok content, their definitions, and example hashtags.
Substance Use Hashtags Network
Community Detection: We first applied established community detection algorithms including Clauset-Newman-Moore (Clauset, Newman, and Moore 2004) and Louvain method (Blondel et al. 2008), but these approaches did not yield clusters that meaningfully captured substance-related discussions. Therefore, we used a hybrid approach combining manual and automated methods to identify thematic groups within the hashtags. We first manually annotated a sample of 251 hashtags and iteratively established thematic categories. As a result, we identified 16 categories representing topics related to substance use and consumption, identity, and community, health and emotion as well as awareness and advocacy. In addition, we created a Misc category for general hashtags, not related to substance use content. The final set of categories and their definitions are shown in Table 1. These categories were then used to prompt GPT-4o (OpenAI 2024) to classify hashtags. When evaluated against human-labeled data, our best-performing prompt achieved an accuracy of $89.36\%$, demonstrating strong reliability in the thematic categorization. We then used this prompt to classify the full set of 2,333 hashtags. The resulting labeled hashtags were mapped back onto the network, creating themed communities that reflect different aspects of substance use discourse on TikTok.
Network Analysis: For both the overall network and each community, we calculated standard network centrality measures including betweenness (Freeman 1977), closeness (Sabidussi 1966), degree (Freeman 1978), and eigenvector centralities (Bonacich 1972). These measures help identify influential hashtags and understand the structure of discourse within and across themes. Additionally, we computed network-level metrics such as average degree, clustering coefficient, and average shortest path length to characterize the overall connectivity patterns and community structure of substance use-related discourse on TikTok.
Characterizing Video Content
Codebook Development: To establish the codebook for coding substance-related TikTok videos, two researchers independently open-coded 160 randomly selected videos. Through this initial coding process, researchers identified nine distinct video topics and coded for substance presence and usage. The physical presence of drugs and related substances was of interest to us in determining which communities and topics were more engaged in depicting substances. The intercoder reliability was strong for coding substance visibility ($\kappa=0.87$) but only fair for video topics ($\kappa=0.35$). After discussing initial disagreements, the topics were refined from nine to seven distinct categories. The coders then had an iterative process for discussing the codes with all researchers on the team to identify coherent topics of videos within the TikTok dataset, resolve any disagreements, and finalize the codebook. The final codebook captured two key dimensions: video topic and substance presence, as shown in Table 3.
Video Selection: We started by selecting all videos that contained at least one of the 2,333 core hashtags, resulting in a dataset of 39,509 videos. To refine this dataset for detailed coding, we identified the top five hashtags per community based on their degree centrality.For each of the 80 selected hashtags, we then filtered videos using a stepwise approach:
  1. 1.
    50 initial videos were acquired from our dataset for each of the 80 hashtags. Ten of each were acquired from each quintile of viewership for the hashtag, achieving representation regardless of virality.
  2. 2.
    Each video was evaluated against three key criteria:
    • Availability: The video is accessible and not moderated or removed by TikTok.
    • Language: The video features English as the primary spoken or written language.
    • Substance Relevance: The video clearly depicts substance-related content, whether through visual cues, discussion, or behaviors.
We continued this process until we identified five videos per hashtag, one per each quintile of engagement, that met all three criteria. In total 977 links were coded, with 852 videos available (87.2%), of those 695 videos were in English (81.5%), and of those 351 videos were substance-related (50.5%). We conducted an inductive content analysis (Elo and Kyngäs 2008) on these 351 selected videos based on the codebook to ensure our final dataset was representative of the substance-use discourse across different communities (See Table 2).
Findings
Filtered Coded
N Videos 39,509 351
Avg. Likes 66.5K 30.9K
Avg. Comments 702.6 305.9
Avg. Shares 3513.2 2652.5
Min. Date 2015-09-22 2019-11-13
Max. Date 2024-05-13 2024-05-09
Table 2: Video statistics of the dataset filtered for RQ1 and the sub-set used for coding in RQ2.
RQ1: Network Structure and Communities
Analysis of hashtags network showed complex, interconnected structures with distinct communities. The overall network contained 2,333 nodes (hashtags) and 46,990 edges (Table C in Appendix), with an average degree of 40.28 and a weighted degree of 513.84 as listed in Table C. These metrics indicate substantial interconnection - each hashtag co-occurs with approximately 40 other hashtags, and when hashtags do co-occur, they appear together frequently (over 500 times on average). The network exhibits small-world properties characterized by a high average clustering coefficient (0.52) coupled with a short average path length (2.20) (Watts and Strogatz 1998). This structure suggests that while substance-related discussions form tight communities, information can still flow efficiently across the network through just two intermediate hashtags on average.
Network centrality analyses revealed two distinct categories of influential nodes. The first category consists of substance-specific hashtags like “#addiction”, “#recovery”, “#sober”, “#wedorecover”, and “#harmreduction”, which scored highly across all centrality measures (betweenness, closeness, degree, and eigenvector). These hashtags function as critical bridges between different substance-related discussions, facilitating information flow between harm reduction, recovery, and other communities. The second category includes general-purpose hashtags like “#fyp” and “#mentalhealth”, which exhibit high centrality but serve primarily to connect substance-related discourse to broader TikTok conversations. This integration with platform-wide hashtags means substance-related content can potentially reach users not explicitly seeking such content through TikTok’s recommendation system.
Our community detection analysis identified 16 different themes (excluding Misc, see Table 1), each revealing distinct patterns in how substance-related content is organized on TikTok (Figure 6 in Appendix). The largest community in this network represents Awareness and Advocacy, containing approximately 12% of all hashtags (280 nodes) and featuring highly central nodes focused on harm reduction (e.g., “#harmreduction”, “#harmreductionworks”), recovery and sobriety (e.g., “#wedorecover”, “#sober”), and public health awareness (e.g., #opioidawareness, #opioidcrisis). The Emotions and Feelings cluster, while smaller with 180 nodes, shows high internal connectivity (clustering coefficient of 0.61) and frequently bridges to other communities through hashtags like “#mentalhealth” and “#anxiety,” suggesting the important role of emotional support in substance-related discussions. The network also revealed unique patterns in how substance-specific discussions are organized. Communities such as Cannabis (142 nodes) and Tobacco/Nicotine (138 nodes), exhibit strong internal clustering but limited external connections. This isolation suggests the discussions stay within their topical boundaries. The Commonly-Misused Substances cluster, though relatively small (120 nodes), contains hashtags often straddling the line between awareness and potential promotion of substance use (e.g., ”#studydrugs”, “#pingertok”, “#smartdrugs”). The Platform-specific cluster (165 nodes) showed distinct characteristics (e.g., “#fyp”, “#viral”), with the highest average betweenness centrality, indicating its crucial role in amplifying substance-related content to broader audiences.
Figure 3: Engagement metrics across select popular hashtags in the coded set. Note, #harmreductionsaveslives has been shorted to HSL.
RQ2: Topics and Characteristics of Videos
Analysis of 351 videos based on our codebook (Table 3) showed that the most prevalent type of content in the videos focused on advocacy and recovery support (33.9%, 119 videos), where creators shared personal recovery journeys, celebrated sobriety milestones, or provided support to others struggling with substance use. For example, one video shows a creator celebrating “100 days sober milestone and encouraging others to seek help.” Satirical or humorous content formed the second largest category (28.2%, 99 videos), with videos often using comedy to address serious topics around substance use. The content is typically presented in the form of acted-out skits or parodies designed to gain engagement. For instance, one video presented a “comedic skit about stereotypical behavior while intoxicated.”
Informational content, typically from claimed medical or professional perspectives, comprised 13.1% (n=46) videos. These videos often featured “doctors explaining the effects of new medications on brain chemistry” or provided educational content about substance effects and risks. A smaller portion focused purely on social media engagement (3.9%, n=14 videos), with creators recording substance-related content primarily to gain followers or views. A sizable portion of the content is in the promotional category, appearing in 7% (n=24) videos. The majority of them portray a demonstration of a vaping device, Juul or lighter, and a few of them promote nootropics or study drugs. Notably, while discussions of substances were common, direct depiction of substances was less frequent - appearing in 25.9% (n=91) of videos, with only 23 of those videos showing active substance use. This suggests a tendency toward discussing rather than displaying substance use.
Dimension Category Definition N (%) Example
Visible Substance Shown Direct visual depiction of drugs or other substances 91(25.9%),with 23(25%) Actively Using Displaying empty prescription medication bottles on a table beside the subject while they reminisce about prior SUD
Not Shown No direct visual depiction of substances 260 (74.1%) discussing recovery journey without showing any substances
Recovery Advocacy Content promoting or raising awareness about recovery efforts 119 (33.9%) Person sharing their 100 days sober milestone and encouraging others to seek help
Satirical Humorous or satirical content about substance use without meaningful substance-related message 99 (28.2%) Comedic skit about stereotypical behavior while intoxicated
Video Topic Informational Educational content about substances from claimed professional/medical perspective 46 (13.1%) Doctor explaining the effects of a new medication on brain chemistry
Promotional Content meant to promote or sell substance-related products 24 (6.8%) Video advertising nicotine vaping products with links to purchase
Social Documentation Content created primarily for social media engagement, focusing on filming substance use 14 (4.0%) Creator recording themselves taking shots at a party while encouraging viewers to like/follow
Trip Reports First-person accounts of substance use experiences 4 (1.1%) Individual describing their experience with a specific substance and its effects
Other Substance-related content that doesn’t fit other categories 45 (12.8%) News clip about substance policy changes
Table 3: Codebook for qualitative analysis of substance-related TikTok videos (N=351)
Figure 4: Sub-networks of Awareness & Advocacy (Pink) and Health Conditions (Blue) to visualize strong overlap between hashtag usage.
Alignment Between Hashtags and Video Content We also examined whether videos’ actual content aligns with their associated hashtag and communities. While each video was coded for its primary topic, we find that videos typically contain hashtags from multiple communities, revealing complex patterns in how substance-related content is tagged and disseminated.
Advocacy, Recovery, and Support Content Videos coded primarily as advocacy and recovery content (119 videos, 33.9%), the largest thematic group in our manual coding, showed strong alignment with two major hashtag communities (see Table 4):
  • Awareness & Advocacy (99 videos, 57.2%): Recovery-focused videos frequently used hashtags like “#harmreduction” and “#wedorecover”. Notably, while recovery-focused hashtags were more commonly used overall, content tagged with harm reduction messaging (particularly “#harmreductionsaveslives”) generated substantially higher user engagement across all metrics - views, likes, comments, and shares (Figure 3).
  • Health Conditions (100 videos, 56.8%): Videos discussing recovery often included medical or health-focused tags like “#addiction”, “#chronicpain”, and “#pain” (Figure 4)
Content of videos seen in this group are typically centered around active or past efforts to recover from substance abuse, discussed in context of cautionary tales, or implicitly imploring the viewer to regard substance abuse with regard and care due to potential harms.
Entertainment and Platform Engagement Entertainment and humor-focused content (99 videos, 28.2%) as well as videos that documented using a substances for engagement (14 videos, 4%) similarly bridged multiple communities. These videos predominantly contained hashtags from the communities:
  • Platform (82 videos, 40.6% of all coded videos): Entertainment videos heavily utilized platform-specific tags like “#fyp” and “#pingtok”
  • Humor (60 videos, 78.9%): Videos using “#addictionhumor” and similar tags often presented substance use through comedy or satire
  • Substance Effects (35 videos, 53.8%): Videos that were satirical or relatable often documented an effect of substance use (e.g., what it’s like being “#high” or “#stoned”). The subjects in these videos may not necessarily be intoxicated but could be enacting or portraying the effects.
Content in this group still discussed topics of substance use and misuse, however the seeming salient topic was humor. A large proportion of videos in this group were meme-based videos, where the original content was either not drug related but brought into drug discussion with filters, captions, or additional context, or was drug related but leaned towards humor and not object lessons.
Figure 5: Distribution of video topics across three levels of substance visibility (no substances shown, substances shown but not used, and active substance use) in TikTok videos.
Educational and Informational Content Videos coded as informational (46 videos, 13.1%) showed connections to multiple professional and health-focused communities:
  • Health Conditions (20 videos, 11.4%): Medical information and health impacts
  • Awareness & Advocacy (29 videos, 16.8%): Harm reduction education
  • Other Substances (16 videos, 24.6%) :
  • Occupation (10 videos, 19.2%): Content from healthcare workers and other professions
Content in this group was thematically clinical or information focused. Humor was used sparingly if at all. The vast majority of videos in this cluster used detailed captions to go over specific facts regarding drug use or were more long-form and had a health professional or recovery advocate speak at length on the topic. Recovery organizations were seen multiple times in the group, with high production value videos.
While advocacy, entertainment, and educational content showed strong alignment with specific hashtag communities, certain clusters served as contextual bridges across all content types. Hashtags from the Commonly Misused Substances (111 videos, 31.6%), Identity & Community (n = 95 videos, 27.1 %), and Occupation (53 videos, 15.1 %) clusters appeared across all seven content categories, suggesting their role in providing context rather than signaling specific content types. For instance, videos tagged with “#opiates” spanned multiple content categories: healthcare workers explaining naloxone administration (Recovery/Advocacy), recovery nurses humorously discussing their reliance on caffeine (Satirical), and medical professionals explaining medication effects (Informational). Similarly, occupation-related hashtags like “#nurse” or “#doctor” and identity-based tags like “#disabled” or “#veteran” appeared across all content. Consumption Method hashtags like “#vape” or “#smoke” (101 videos, 28.8% of videos) served a more specific bridging function, describing substance use behaviors across content types.
RQ3: Effects of Substance Visibility
Given these patterns in how content creators use hashtags to frame and contextualize their videos, we next examined a key aspect of substance-related content: the inclusion of substances or their related paraphernalia themselves in the videos with respect to context and engagement. This analysis revealed distinct patterns in content distribution across three categories: no substances shown (n=260), substances shown but not used (n=68), and active substance use (n=23). Videos showing active substance use were predominantly focused on documentation for social media (approximately 40% of videos, Figure 5), with a stronger tendency towards lifestyle or vlog-style content. In contrast, recovery advocacy content was most prevalent in videos without substance visibility, while promotional content appeared most frequently in videos that showed but did not depict active use of substances (e.g., videos promoting nicotine products). Within videos showing substances, content predominantly featured alcohol, nicotine, or cannabis products, with only one video collage containing “hard drugs” such as Percocet, Adderall, and Xanax. A Kruskal-Wallis H (Kruskal and Wallis 1952) test was performed to compare the means in engagement metrics (views, likes, comments, and shares) between the substance presence categories. Notably, there were no statistically significant differences between groups. This suggests that user engagement on TikTok is driven more by content type and presentation than by the mere presence or use of substances.
Discussion
Social media’s impact on substance use behaviors has been a growing concern, with platforms often criticized for normalizing or promoting dangerous behaviors (Curtis et al. 2018; Moreno and Whitehill 2014; Lobstein et al. 2017; Rutherford et al. 2023; Romer and Moreno 2017). Our analysis of TikTok, however, revealed a more nuanced reality where recovery and harm reduction messaging dominate substance-related discourse. We found evidence of TikTok emerging as a significant platform for recovery support while maintaining its core identity as an entertainment platform.
The Power of Digital Recovery Narratives The most striking finding is the prevalence of recovery-focused content, which typically feature first-person narratives celebrating sobriety milestones or sharing recovery journeys. As Slade et al. (2021) note, such “digital recovery narratives” can increase knowledge of evidence-based treatment options, normalize addiction experiences, and encourage help-seeking behaviors. The impact of these narratives appears to be bidirectional - benefiting creators and viewers. For creators, sharing recovery stories can reinforce recovery-focused, collective identities and provide accountability through community support (Beck et al. 2023). For viewers, these narratives offer realistic models of recovery success, reduce feelings of isolation, and demonstrate that recovery is possible across different circumstances (Jacobson 2001; Kaiser et al. 2020; Corvini, Golomski, and Burns 2024).
Balancing Entertainment and Support in Health Communication The significant presence of humorous content (28.2%) alongside serious recovery narratives reflects TikTok’s unique ability to facilitate health communication through multiple approaches. Rather than undermining serious messages, humor appears to serve as an additional entry point for engaging with recovery-related content. The use of humor in substance-related content may serve important psychological functions, as research has shown that comedy can help individuals process difficult experiences (Martin 2002, 2003) and reduce stigma around mental health challenges (Corrigan et al. 2014). In the context of substance use and recovery on TikTok, videos that integrate humor appear to make the challenges around substance use more approachable while facilitating community connection through shared experiences. This balance between entertainment and support may be particularly effective for reaching younger audiences. However, this dual nature also presents challenges. While most humorous content avoids glorifying substance use some videos may inadvertently normalize concerning behaviors through comedy. This tension between engagement and responsible messaging represents a key challenge for platforms hosting health-related content.
Network Structure Facilitates Community Building The network analysis revealed a highly interconnected structure with small-world properties, enabling both focused discussions within communities and efficient information flow between them. Hashtags serve multiple functions: some act as content summarizers (e.g., “#harmreduction”, “#recovery”), while others serve as contextualizers (e.g., identity-based tags, occupation tags). Within this network structure, we observed the presence of algospeak in the form of coded or evasive hashtags (e.g., “#drvg”, “#drvgzx”, “#dr00gtiktok”) that are designed to evade content moderation. Our content analysis revealed that these coded hashtags were predominantly associated with satirical and relatable content, with a small subset appearing in recovery advocacy videos. This pattern suggests these alternative spellings may serve less to promote substance use and more to facilitate humor-based discussions and shared experiences around substance use, though this practice could complicate platforms’ efforts to effectively moderate substance-related content (Klug, Steen, and Yurechko 2023; Levine 2022). This dual functionality helps explain why content tagged with harm reduction and recovery-focused hashtags received substantially higher engagement despite appearing less frequently in our sample.
Looking Forward: Implications for Social Media Health Communication
Our findings have important implications for how social media platforms approach substance-related content, not just on TikTok but across all platforms featuring short-form video content like Instagram Reels and YouTube Shorts. The success of recovery narratives suggests that brief, personal videos may be particularly effective for health communication. Rather than focusing solely on restricting content, platforms might consider ways to actively promote recovery and harm reduction content, given its demonstrated appeal to users.
The emergence of TikTok as a recovery support space (Russell et al. 2022) offers valuable lessons for content moderation and platform design. While current approaches appear to effectively limit harmful content while allowing space for support and advocacy, refinements could address the challenges of moderating humor-based content. As short-form video becomes an increasingly dominant mode of social media communication, particularly among younger users, understanding how to foster supportive recovery communities while mitigating potential risks becomes important. Several areas warrant further research. While we found evidence of TikTok serving as a positive space for recovery support, there is a pressing need to understand the real-world impacts of this online discourse. Future studies should examine how exposure to recovery content affects help-seeking behaviors and treatment outcomes. Additionally, researchers should investigate potential risks and benefits for people who use drugs who engage with or create TikTok content. This is particularly important for content creators who publicly identify as being in active recovery, as sharing personal narratives may have complex psychological and social implications that are not yet well understood (Rennick-Egglestone et al. ). Additionally, investigating the intersection of online and offline support networks could provide insights into how digital recovery communities complement traditional recovery support services, optimizing online and in-person interventions. This work would be particularly valuable for reaching younger populations, who may be more likely to seek initial support through social media platforms.
Conclusion
This study explored the use of over 2.3K substance-related hashtags used on the popular short-form social media platform, TikTok. Through a semi-automatic community detection process, we identified 16 distinct hashtag communities, including Advocacy and Awareness, Health Conditions, and Humor. Our social network analysis revealed a highly interconnected small-world network (clustering coefficient 0.52, average path length 2.20) with influential hashtags like “#addiction”, “#recovery”, and “#sober” serving as bridges between communities. Through manual coding of 351 representative videos, we found that Recovery Advocacy content dominated (33.9% of videos), followed by Satirical content (28.2%), suggesting TikTok serves primarily as a platform for support rather than substance promotion. This is further supported by our finding that substances were present in only 91 videos (26%), with active use shown in just 23 videos. The alignment between hashtag communities and video content, particularly in recovery-focused videos, indicates organic community formation rather than attempts to evade content moderation. These findings have important implications for platform policies and interventions - while the prevalence of recovery-focused content suggests TikTok may serve as a valuable space for support and advocacy, the significant presence of humorous content alongside serious advocacy highlights the need for nuanced content moderation approaches. Future research should examine how these digital recovery communities complement traditional recovery support services and their real-world impacts on help-seeking behaviors and treatment outcomes, particularly among youth.
Limitations
Several limitations of our study warrant discussion. First, our findings may not fully represent the entire TikTok ecosystem, as they are influenced by limitations in data availability and collection processes. TikTok’s (unofficial) API serves content through its recommendation algorithm, which means we received videos that did not necessarily contain our queried hashtags. While we attempted to address this by sampling videos across engagement quartiles to ensure representativeness, the algorithm’s opacity makes it difficult to guarantee complete coverage of substance-related content. This algorithmic bias may have also led to over-representation of certain creators whose content is favored by TikTok’s recommendation system. TikTok, in our experience, also cuts off data requests arbitrarily. Our approach of using 1,000 videos per hashtag was due to an alignment with the average number of videos we could reliably retrieve before being cut off on that request. Second, the dynamic nature of social media data presented challenges - some videos included in our initial network analysis were later unavailable for content coding due to accounts being banned or made private, a particular issue in communities discussing sensitive topics like substance use. Third, methodological challenges arose around identifying substance-related content, as users often employ common words as coded references to substances, making it difficult to definitively categorize some content as substance-related versus general discussion. This challenge was compounded by the platform’s rapid evolution of slang and terminology. Additionally, some videos contained duplicate hashtags, which could affect network metrics, and less popular hashtags may be underrepresented in our sample despite efforts to ensure broad coverage. These limitations highlight the challenges of studying substance-related content on social media platforms and suggest the need for complementary research approaches to fully understand these communities.
Appendix A Appendix
Appendix B Full Hashtag Network with Communities
Figure 6 shows a visualization of the complete hashtag co-occurrence network, containing 2,333 nodes (hashtags) and 46,990 edges. The nodes are colored according to their community membership as determined by our semi-automated community detection process. The network layout was generated using a Fruchterman-Reingold algorithm, which positions nodes with stronger connections closer together. The central position of recovery-focused hashtags like “#addiction”, “#recovery”, and “#harmreduction” is clearly visible, demonstrating their role as bridges between communities. Node sizes are scaled according to their degree centrality, highlighting the most connected hashtags in the network. This visualization helps illustrate both the small-world properties of the network (clustering coefficient 0.52, average path length 2.20) and the organic formation of thematic communities around substance-related discourse on TikTok.
Figure 6: The overall network of 2,333 hashtags related to substance use with colors representing each community as shown in the legend
Appendix C Network Centrality Measures per Community
Table C presents detailed network statistics and centrality measures for both the complete network and each of the 17 identified communities including Misc. For each community, we report basic network metrics (number of nodes, edges, average degree, weighted degree, clustering coefficient, and average shortest path length) as well as the top 5 most central nodes according to four different centrality measures (betweenness, closeness, degree, and eigenvector centrality). The whole network shows high connectivity (average degree 40.28) and clustering (0.52), with core hashtags like “addiction”, “recovery”, and “sober” consistently ranking highly across all centrality measures.
Communities
Documentation of
Use for
Social Media
Informational Content Other Promotional Content Recovery Advocacy
Satirical and
Relatable Content
Trip Reports Total
Platform 8 (4.0%) 20 (9.9%) 25 (12.4%) 15 (7.4%) 59 (29.2%) 74 (36.6%) 1 (0.5%) 202
Health Conditions 0 (0.0%) 20 (11.4%) 17 (9.7%) 3 (1.7%) 100 (56.8%) 34 (19.3%) 2 (1.1%) 176
Awareness and Advocacy 1 (0.6%) 29 (16.8%) 12 (6.9%) 1 (0.6%) 99 (57.2%) 29 (16.8%) 2 (1.2%) 173
Misc 4 (3.2%) 8 (6.4%) 24 (19.2%) 7 (5.6%) 42 (33.6%) 39 (31.2%) 1 (0.8%) 125
Commonly Misused Substances 1 (0.9%) 13 (11.8%) 20 (18.2%) 1 (0.9%) 39 (35.5%) 36 (32.7%) 0 (0.0%) 110
Consumption Method 3 (3.0%) 11 (11.1%) 12 (12.1%) 14 (14.1%) 22 (22.2%) 37 (37.4%) 0 (0.0%) 99
Emotions and Feelings 4 (4.1%) 8 (8.2%) 14 (14.3%) 4 (4.1%) 32 (32.7%) 36 (36.7%) 0 (0.0%) 98
Identity and Community 2 (2.1%) 8 (8.5%) 9 (9.6%) 2 (2.1%) 51 (54.3%) 20 (21.3%) 2 (2.1%) 94
Humor 0 (0.0%) 2 (2.6%) 4 (5.3%) 1 (1.3%) 8 (10.5%) 60 (78.9%) 1 (1.3%) 76
Other Substances 2 (3.1%) 16 (24.6%) 11 (16.9%) 0 (0.0%) 21 (32.3%) 15 (23.1%) 0 (0.0%) 65
Substance Effects 7 (10.8%) 4 (6.2%) 13 (20.0%) 4 (6.2%) 6 (9.2%) 28 (43.1%) 3 (4.6%) 65
Location 3 (5.0%) 4 (6.7%) 13 (21.7%) 2 (3.3%) 9 (15.0%) 28 (46.7%) 1 (1.7%) 60
Alcohol 8 (13.6%) 6 (10.2%) 6 (10.2%) 0 (0.0%) 19 (32.2%) 18 (30.5%) 2 (3.4%) 59
Tobacco/Nicotine 2 (3.4%) 9 (15.3%) 5 (8.5%) 12 (20.3%) 16 (27.1%) 15 (25.4%) 0 (0.0%) 59
Occupation 1 (1.9%) 10 (19.2%) 9 (17.3%) 1 (1.9%) 22 (42.3%) 9 (17.3%) 0 (0.0%) 52
Cannabis 0 (0.0%) 3 (8.1%) 8 (21.6%) 0 (0.0%) 4 (10.8%) 21 (56.8%) 1 (2.7%) 37
Cognitive Enhancement 1 (4.0%) 7 (28.0%) 3 (12.0%) 7 (28.0%) 4 (16.0%) 3 (12.0%) 0 (0.0%) 25
Table 4: Distribution of communities corresponding with each hashtag used in the 351 analyzed videos per content topic.
Network #N #E Avg D Avg WD Avg C Avg SP Betweenness Closeness Degree Eigenvector
Whole Network 2333 46990 40.28 513.84 0.52 2.20 addiction, recovery, sober, wedorecover, harmreduction addiction, recovery, sober, wedorecover, harmreduction addiction, recovery, sober, wedorecover, harmreduction addiction, recovery, sober, wedorecover, harmreduction
Awareness & Advocacy 279 2164 15.51 334.30 0.89 1.98 harmreduction, wedorecover, harmreductionsaveslives, harmreductionworks, sober harmreduction, harmreductionsaveslives, wedorecover, harmreductionworks, sober harmreduction, harmreductionsaveslives, wedorecover, harmreductionworks, sober harmreduction, harmreductionsaveslives, wedorecover, harmreductionworks, narcansaveslives
Health Conditions 277 1522 10.99 224.09 0.85 2.02 addiction, recovery, sobriety, pain, chronicpain addiction, recovery, sobriety, pain, chronicpain addiction, recovery, sobriety, pain, chronicpain addiction, recovery, sobriety, pain, detox
Emotions & Feelings 172 305 3.55 20.65 0.31 2.39 peace, onedayatatime, friend, dream, serenity peace, onedayatatime, dream, serenity, friend peace, onedayatatime, dream, friend, serenity peace, onedayatatime, dream, serenity, friend
Platform 183 146 1.60 39.58 0.12 2.44 pingtok, dr00gtiktok, pingertok, chat, fyp pingtok, dr00gtiktok, pingertok, fyp, fyp:) pingtok, dr00gtiktok, pingertok, chat, stonertokfyp pingtok, dr00gtiktok, pingertok, chat, fyp
Identity & Community 155 97 1.25 8.22 0.09 3.21 cleanandsober, na, pwud, 30s, women pwud, na, cleanandsober, women, lgbt cleanandsober, pwud, na, 30s, mujeres cleanandsober, pwud, na, lgbtqia, calisober
Commonly Misused Substances 152 553 7.28 60.47 0.38 2.47 opioids, opiates, coke, molly, fent opioids, coke, opiates, molly, uppers opioids, opiates, coke, molly, fent opioids, coke, opiates, fent, molly
Location 116 19 0.33 1.26 0.00 1.80 chinatown, paradise, yucatan, arnolds, filmteyvatislands chinatown, paradise, india, china, texas chinatown, paradise, yucatan, arnolds, india chinatown, india, china, texas, egypt
Cannabis 51 229 8.98 39.25 0.55 2.04 maryjane, weed, cannabis, skunk, herb maryjane, weed, cannabis, herb, grass maryjane, weed, cannabis, herb, grass maryjane, cannabis, weed, herb, grass
Alcohol 43 92 4.28 56.19 0.41 1.79 drunk, tipsy, hammered, shitfaced, bars drunk, hammered, tipsy, shitfaced, bars drunk, hammered, tipsy, shitfaced, bars drunk, hammered, tipsy, bars, shitfaced
Cognitive Enhancement 24 38 3.17 45.00 0.46 1.64 nootropics, modafinil, smartdrugs, smartpills, focus nootropics, smartdrugs, modafinil, smartpills, focus nootropics, smartdrugs, modafinil, smartpills, focus nootropics, smartdrugs, modafinil, smartpills, focus
Substance Effects 41 33 1.61 23.76 0.16 2.42 high, wasted, energy, blackout, clarity high, wasted, lit, crash, shadowpeople high, wasted, blackout, lit, downers high, wasted, lit, downers, crash
Consumption Method 41 62 3.02 89.80 0.30 1.86 smoke, vape, smoking, vaper, vaping smoke, vaper, vape, vaping, vaped smoke, vaper, vape, vaping, vaped smoke, vaper, vaping, vaped, vape
Tobacco Nicotine 31 115 7.42 68.90 0.70 1.70 ecigs, ecig, cigarette, nicotine, juul ecigs, cigarette, ecig, nicotine, juul ecigs, cigarette, ecig, nicotine, juul cigarette, ecig, ecigs, nicotine, juul
[…]
Occupation 73 5 0.14 0.93 0.00 1.50 copilots, truckdrivers, actress, nursing, outreachworker copilots, truckdrivers, aviation, construction, creator copilots, truckdrivers, aviation, work, construction copilots, aviation, construction, creator, truckdrivers
Misc 565 2367 8.38 139.01 0.29 2.46 cat, fire, rock, g, pink cat, fire, g, rock, pink cat, fire, g, rock, x cat, fire, g, x, rock
Table 5: Centrality measures for the whole network and each community. #N=Number of Nodes, #E=Number of Edges, Avg D=Average Degree, Avg WD=Average Weighted Degree, Avg C=Average Clustering, Avg SP=Average Shortest Path. The top five nodes are presented per centrality measure.
Appendix D Hashtag Community Distribution per Content Topic
Table 4 presents a cross-tabulation of hashtag communities and video content topics, allowing us to examine how different communities align with various types of content. This analysis reveals notable patterns in how hashtags from different communities are employed across content types. Recovery Advocacy content, the most prevalent in our dataset, shows strong representation across multiple hashtag communities, particularly drawing from the Awareness & Advocacy and Health Conditions communities. Satirical content, our second most common category, shows interesting patterns of hashtag usage, frequently incorporating hashtags from Platform and Humor communities while also drawing from more serious communities - suggesting complex layering of messaging in humorous content. Informational content shows strong alignment with Professional and Health Condition hashtags, while Promotional content primarily uses hashtags from Consumption Method and Tobacco/Nicotine communities. This distribution helps validate our community detection approach by demonstrating meaningful alignment between hashtag communities and content types, while also revealing how creators strategically combine hashtags from multiple communities to reach diverse audiences.